# What Makes for Good Image Captions?

### v2

@@ C.2 Caption Pyramids
For the caption merging function $\sigma_{\mathrm{merge}}$, we adopt and compare a variety of Large Language Models (LLMs) as its implementation, including the Gemma family (2B and 7B versions), LLaMA2 family (7B chat and 13B chat), Qwen-1.5-7B Chat, Mistral 7B, and a mixture-of-expert model Mixtral 8x7B. The Mixtral 8x7B model has capabilities similar to ChatGPT-3.5 and is one of the top-performing open-source LLMs. All inference is performed in FP16 precision, except for the large Mixtral 8x7B model, for which we use 8-bit quantization to fit it into a single NVIDIA A800 GPU. The Mixtral 8x7B is used as the default LLM for caption merging.
[…]
@@
We employ the prompt shown in Table 2 for caption merging, where the "Assistant Generation Prefix" is injected after the instructions to control the model output format. All generation parameters are set to the default values provided by the source repository. For splitting function $\sigma_{\mathrm{merge}}$, we adopt the most straightforward implementation by splitting the input image into four equal-sized patches.
C.3 Human annotation baseline
@@
In Fig. 2, the human annotation for short captions represents the accuracy of a single-sentence caption drawn from the MS-COCO annotation, while the human annotation for detailed captions refers to five MS-COCO caption annotations concatenated with the prefix "The following are several captions of this image written by different people: " added to the front.
Appendix D Additional Experiments and Further Analysis
D.1 Image Paragraph Captioning
@@
The image paragraph captioning dataset contains human-annotated single-paragraph descriptions for Visual Genome images. We use its testing split, which consists of 2,492 samples. We employ both the reference-free metric CLIPScore [62] and the reference-based metric METEOR [63] to evaluate the quality of the captions.
@@
CLIPScore measures the similarity between the image and text features extracted by the CLIP model (we use the standard OpenAI pretrained ViT-Base-32). The underlying assumption is that CLIP encoders are capable of extracting semantic information and can represent the importance score $A$; thus, a higher CLIPScore correlates with higher information sufficiency.
@@
The reference-based metric METEOR is widely adopted for evaluations in image captioning and natural language generation (e.g., machine translation). It measures the word-level similarity between model-generated captions and human-generated captions. The underlying assumption is that human annotations optimize the information sufficiency objective, so if a model behaves similarly to human annotations, it achieves high information sufficiency.
@@
The results are shown in Table 3, where we also list the performance of previous fully-supervised models and few-shot models. Once again, our PoCa method provides information sufficiency improvement according to both the reference-free metric CLIPScore and the reference-based metric METEOR across all three families of LVLMs. These results further demonstrate the effectiveness of the PoCa method in enhancing the quality and informative content of the generated captions.
@@
Fully Supervised Models Regions-Hierarchical [28] - 15.95
RTT-GAN [64] - 17.12
HSGED [65] - 18.33
SCST [66] - 17.86
CAE-LSTM [67] - 18.82
Few-shot & Zero-shot Models BLIP-2 3-shot - 10.8
@@
D.2 Caption Merging Strategies
In this section, we compare the effectiveness of different implementations of the merging function $\sigma_{\mathrm{merge}}$. First, we compare various LLMs introduced in Section C.2. As shown in Table 4, compared to the global caption baseline, every LLM yields performance improvement, except for the smallest Gemma-2B-IT model. We also provide an ablation on prompting, where we replace the default prompt shown in Table 2 with a naive prompt of "merge these captions". This ablation results in a slight decrease in accuracy and a significant increase in caption length, which further violates the minimal redundancy objective.
@@ D.2 Caption Merging Strategies
D.3 Analysis of VQA-based Caption Evaluation
Prompt for LLM-based VQA Evaluation System Message: You will be given a caption of an image, and your task is to try to answer the question based on the caption. If the relevant information is not present in the caption, try your best to guess the answer. You shouldn’t provide any rationale or explaination in your response, just give the answer only. The answer can be a number, a single word or a short phrase, plese make your response as short, simple and clear as possible. User: Image Caption: {image caption} Question: {question} Assistant Generation Prefix: The most possible answer is:
Table 5: Prompt for LLM-based VQA Evaluation.
@@ D.3 Analysis of VQA-based Caption Evaluation
In the VQA-based evaluation, we adopt text-only LLMs for VQA inference with image captioning input to assess caption quality. This section provides a detailed analysis of this approach. Using the instruction given in Table 5, we prompt different LLMs to generate answers and evaluate the accuracy based on exact matching and Natural Language Inference (NLI) based evaluation. The NLI evaluation classifies a pair of statements, "The answer to this question is {ground truth}" and "The answer to this question is {generated answer}", into entailment, neutral, and contradiction, where entailment outputs are regarded as successful. Compared to exact matching, NLI evaluation measures the correctness of answers at a semantic level and can tolerate low-level differences.
@@
Limitations
VQA Evaluation. While the VQA-based evaluation provides a useful measure of caption quality in terms of information sufficiency, it has limitations. The questions used for evaluation may not comprehensively cover all of the important semantic units, resulting a sub-optimal estimation of the importance score $A$. In addition, due to resource constraints, we use a 5,000 question subset from the full VQAv2 dataset. To test its reliability, we run default caption generation with 5 models, together with human annotated caption, resulting in a total of (5+1)$\times$2=12 data points combining short and long captioierns. The Pearson correlation coefficient between 5k subset accuracy and full dataset accuracy is 0.8519 – although already quite high, it still introduce some degree of noise for model performance evaluation.
Computational Efficiency. Our implementation of PoCa involves more inferences to generate captions and prompting LLM for fusing the local and global captions. These multiple inference steps and the use of large models can lead to increased computational costs. This computational overhead may be a concern, especially in resource-constrained environments or when processing a large number of images. One potential solution is to finetune an image captioning model on the captions generated by PoCa. By doing so, the knowledge captured by PoCa can be distilled into the finetuned model, allowing for a single inference pass during deployment, while still benefiting from the enhanced caption quality achieved by PoCa. Similar approach of knowledge distillation has been adopted in other literature, such as [33, 69, 70].
Broader Impact
The proposed PoCa method has the potential to positively impact various applications, such as improving accessibility for visually impaired individuals and enhancing information retrieval. However, it is crucial to consider and address potential ethical implications, including biases inherited from LLMs, privacy concerns related to revealing sensitive information, and the environmental impact of training and deploying large-scale models. Responsible deployment, clear communication of capabilities and limitations, and ongoing efforts to mitigate biases and ensure fairness are essential for developing image captioning systems that are not only informative and accurate but also socially responsible and inclusive.
Appendix E PoCa Examples
We show some example of PoCa caption merging, where the images are sourced from the test split of the image paragraph captioning dataset. Captions are generated by LLaVA-1.5-7B model and merged by Mixtral-8x7B LLM. It shows that local captions successfully captured some errors that are hallucinated or omitted by the global caption, while PoCa successfully aggregated such information (marked by green) and recompile it into a human-interpretable caption.
@@ Appendix E PoCa Examples
Global Caption The image features a group of four people dressed in old-fashioned
clothing, standing together in a train station. They are posing for a picture in front of a train,
which is located on the left side of the scene. The people are dressed in various styles, with some
wearing ties and hats. There are two handbags visible in the scene, one near the center and the
other towards the right side. The overall atmosphere of the image suggests a vintage or historical
setting.
Top-left: The image features a man standing in front of a train, which is prominently
displayed in the scene. The train is positioned towards the right side of the image, and the man is
standing closer to the left side. The man appears to be posing for a picture in front of the train.
In the background, there is a building with a clock on its side, adding to the overall atmosphere of
the scene. Additionally, there are two other people in the image, one located near the left edge and
the other towards the right side.
Bottom-left: The image features a train station with a train on the tracks. There are
several people in the scene, some of them standing near the train, while others are scattered
throughout the station. A man is standing close to the train, possibly waiting for it to depart. In
addition to the people, there are a few handbags visible in the scene, with one near the center of
the image and another closer to the right side. The presence of these handbags suggests that the
people in the station might be travelers or commuters.
Top-right: The image features a group of three people dressed in top hats and
Victorian-style clothing, standing together in a large room. They appear to be posing for a picture,
possibly in a train station or a similar setting. The people are positioned in the center of the
room, with one person on the left, another in the middle, and the third on the right side. The room
has a few chairs placed around the area, with one chair located near the left side of the room,
another in the middle, and a third chair on the right side. The overall atmosphere of the scene
suggests a historical or themed event.
Bottom-right: The image features a man and a woman standing next to each other, both
dressed in old-fashioned clothing. The woman is wearing a green dress, while the man is dressed in a
suit with a tie. They appear to be posing for a picture, possibly in a historical setting. In the
scene, there are also two handbags visible, one near the woman and the other closer to the man. The
overall atmosphere of the image suggests a vintage or historical theme.
PoCa Merged Caption: A group of four people in old-fashioned clothing, including two in
top hats, pose for a picture in a train station with a vintage atmosphere. A train is visible on the
left, and two handbags are spotted near the center and right. Two other people are scattered
throughout the scene, with one man standing close to the train.
Human Annotation: Four people wearing old-fashioned clothing stand posing in front of the
camera. They are in a large building with a bright ceiling, and there is a black train behind them.
There are two men on either side of two women. The man on the far left has a black suit, black
shoes, and a black top hat. The woman to his right is wearing a white, pink, and green dress, with a
white hat, and a broach pinned to her chest. She is holding a white satchel in her white-gloved
hands. The woman next to her is also wearing white gloves and holding a white satchel. This woman
has a green dress on, with lace on the trim. The man next to her is wearing a checkered vest
underneath a dark long coat and dark pants. He has a black top hat.
@@
Global Caption The image depicts a group of people gathered around a dining table in a
restaurant. They are enjoying a meal together, with several pizzas placed on the table. There are
multiple cups and bottles scattered around the table, indicating that they are drinking beverages as
well. The people are seated on various chairs placed around the table. Some of them are engaged in
conversation, while others are focused on their meal. The atmosphere appears to be relaxed and
social, as the group shares a meal and spends time together.
Top-left: The image depicts a group of people gathered in a room, possibly a restaurant or
a bar. There are at least nine people visible in the scene, with some sitting at a dining table and
others standing around the room. The room features a spiral staircase, which adds a unique
architectural element to the space. A few chairs are placed around the dining table, and a cup can
be seen on the table. There is also a bowl and a spoon placed nearby. The atmosphere appears to be
lively and social, with people enjoying their time together.
Bottom-left: The image features a group of people sitting around a dining table, enjoying
a meal together. There are at least five people visible in the scene, with some sitting closer to
the table and others further away. The table is set with various food items, including a pizza, a
bowl, and a sandwich. There are also multiple cups and a cell phone placed on the table. A bottle
can be seen on the table as well. The dining table occupies a significant portion of the image, with
the people sitting around it, creating a warm and inviting atmosphere.
Top-right: The image features a man with a long beard sitting at a dining table in a
restaurant. He is looking at the camera, possibly posing for a picture. The man is wearing a blue
shirt and appears to be enjoying his time at the restaurant. The dining area is furnished with
multiple chairs and dining tables. There are two chairs visible in the scene, one near the man and
another further away. Two dining tables can be seen, one in the foreground and another in the
background. In the background, there is a TV mounted on the wall, likely providing entertainment for
the restaurant’s guests.
Bottom-right: The image features a person sitting at a dining table with a plate of pizza
in front of them. The table is set with a glass of beer, a cup, and a pitcher of water. There are
also two additional cups on the table. The person is holding a fork, ready to enjoy their meal. The
dining table is surrounded by chairs, with one chair on the left side and another on the right side.
The scene appears to be a casual dining experience, with the person enjoying their pizza and beer.
PoCa Merged Caption: A group of people, including a man with a beard, enjoy a meal in a
restaurant with various pizzas, sandwiches, and beverages on a table. Nine people are gathered in a
room with a spiral staircase and chairs around the dining table. The atmosphere is lively and
social. (169 characters)
Human Annotation: three men are talking. they all have bears on their face. there is a
blue shirt on the man. the shirt has flowers on it.

### v3

@@ 5.1 Implementation Details
Caption Pyramids. For the caption merging function $\sigma_{\mathrm{merge}}$, we adopt and compare a variety of Large Language Models (LLMs) as its implementation, including the Gemma family (2B and 7B versions), LLaMA2 family (7B chat and 13B chat), Qwen-1.5-7B Chat, Mistral 7B, and a mixture-of-expert model Mixtral 8x7B. The Mixtral 8x7B model has capabilities similar to ChatGPT-3.5 and is one of the top-performing open-source LLMs. All inference is performed in FP16 precision, except for the large Mixtral 8x7B model, for which we use 8-bit quantization to fit it into a single NVIDIA A800 GPU. The Mixtral 8x7B is used as the default LLM for caption merging.
@@
We employ the prompt shown in Table 5 for caption merging, where the "Assistant Generation Prefix" is injected after the instructions to control the model output format. All generation parameters are set to the default values provided by the source repository. For splitting function $\sigma_{\mathrm{merge}}$, we adopt the most straightforward implementation by splitting input image into four equal-sized patches.
5.2 PoCa Improves Task Sufficiency
Prompt for LLM-based VQA Evaluation
System Message:
You will be given a caption of an image, and your task is to try to answer the question based on the caption. If the relevant information is not present in the caption, try your best to guess the answer. You shouldn’t provide any rationale or explaination in your response, just give the answer only. The answer can be a number, a single word or a short phrase, plese make your response as short, simple and clear as possible.
User:
Image Caption: {image caption}
Question: {question}
Assistant Generation Prefix:
The most possible answer is:
Table 1: Prompt for LLM-based VQA Evaluation.
5.2.1 VQA-based Evaluation
In this section, we conduct quantitative evaluations to study whether the PoCa method is able to improve the task sufficiency term in the objective (Definition 3). We adopt the VQA-v2 Goyal et al. (2017) dataset, which is built upon the MS-COCO Chen et al. (2015) dataset and contains multiple questions per image.
We employ a text-only LLM to answer the questions in VQA-v2 based on the generated captions, using the instruction shown in Table 1. The questions serve as a proxy for the importance score $A$ in Definition 3, and the accuracy of the LLM-generated correct answers becomes an estimation of the task sufficiency term. We use LLaMA2-Chat-13B for VQA inference; a detailed analysis of different LLMs for VQA evaluation will be provided in Section B.
Figure 3: VQA accuracy using captions generated by different image captioning models and the proposed PoCa method. The PoCa method consistently improves the VQA performance across all models and caption types.
@@ 5.2.1 VQA-based Evaluation
Figure 3 provides the evaluation results on 5,000 questions in the VQA-v2 validation split. The human annotation for short captions represents the accuracy of a single-sentence caption drawn from the MS-COCO annotation, while the human annotation for detailed captions refers to five MS-COCO caption annotations concatenated with the prefix "The following are several captions of this image written by different people: " added to the front.
As can be seen, our proposed PoCa method (green) consistently yields performance gains across all three examined LVLMs. The scale of improvement ranges from 0.48% (MobileVLM-v2 detailed captions) to 2.10% (LLaVA-1.5 short captions). Interestingly, we find that detailed captions do not necessarily correlate with better information coverage, as the detailed captions generated by MobileVLM-v2 underperform the single-sentence captions generated by InternVL. The comparison also shows that human annotations may not be optimal for certain scenarios, since several groups of LVLM-generated captions can yield higher VQA accuracy compared to that of human annotators.
5.2.2 Image Paragraph Captioning
@@
The image paragraph captioning dataset contains human-annotated single-paragraph descriptions for Visual Genome images. We use its testing split, which consists of 2,492 samples. We employ both the reference-free metric CLIPScore Hessel et al. (2021) and the reference-based metric METEOR Banerjee and Lavie (2005) to evaluate the quality of the captions.
@@
CLIPScore measures the similarity between the image and text features extracted by the CLIP model (we use the standard OpenAI pretrained ViT-Base-32). The underlying assumption is that CLIP encoders are capable of extracting semantic information and can represent the importance score $A$; thus, a higher CLIPScore correlates with higher task sufficiency.
@@
The reference-based metric METEOR is widely adopted for evaluations in image captioning and natural language generation (e.g., machine translation). It measures the word-level similarity between model-generated captions and human-generated captions. The underlying assumption is that human annotations optimize the task sufficiency objective, so if a model behaves similarly to human annotations, it achieves high task sufficiency.
@@
The results are shown in Table 2, where we also list the performance of previous fully-supervised models and few-shot models, including Regions-Hierarchical Krause et al. (2017), RTT-GAN Liang et al. (2017), HSGED Yang et al. (2020), SCST Melas-Kyriazi et al. (2018), CAE-LSTM Wang et al. (2019), VLIS Chung and Yu (2023). Once again, our PoCa method provides task sufficiency improvement according to both the reference-free metric CLIPScore and the reference-based metric METEOR across all three families of LVLMs. These results further demonstrate the effectiveness of the PoCa method in enhancing the quality and informative content of the generated captions.
@@
Fully
Supervised
Models
Regions-Hierarchical - 15.95
RTT-GAN - 17.12
HSGED - 18.33
SCST - 17.86
CAE-LSTM - 18.82
Few-shot &
Zero-shot
Models
BLIP-2 3-shot - 10.8
@@
5.3 Ablation Study and Further Analysis
5.3.1 Caption Merging Strategies
In this section, we compare the effectiveness of different implementations of the merging function $\sigma_{\mathrm{merge}}$. First, we compare various LLMs introduced in Section 5.1. As shown in Table 3, compared to the global caption baseline, every LLM yields performance improvement, except for the smallest Gemma-2B-IT model. We also provide an ablation on prompting, where we replace the default prompt shown in Table 5 with a naive prompt of "merge these captions". This ablation results in a slight decrease in accuracy and a significant increase in caption length, which further violates the minimal redundancy objective.
@@ 5.3.1 Caption Merging Strategies
5.3.2 PoCa Does Not Sacrifices Minimal Redundancy
Previous evaluations primarily focused on the task sufficiency of image captions. However, as our objective also contains other terms, it is crucial to evaluate whether the performance gain brought by PoCa is achieved by significantly sacrificing other objectives. In Table 4, we present the length statistics for captions used in the previous task sufficiency comparison. We calculate the average number of words in default captions and PoCa captions and note their differences in the $\pm\Delta$” column.
The results show that PoCa does not exhibit a significant trend of either increasing or decreasing the length of captions. Among the six comparisons, PoCa compresses the length in four cases and extends the length in two cases. This empirically demonstrates that using LLMs as $\sigma_{\mathrm{merge}}$ in the PoCa model does not significantly violate the minimal redundancy objective $H(\tilde{Y}_{\mathrm{merged}})$.
Table 4: Comparison of caption lengths between default captions and PoCa captions on VQAv2 and Img2P datasets.
LVLM VQAv2 Img2P
Default PoCA $\pm\Delta$ Default PoCA $\pm\Delta$
MobileVLM-v2-1.7B 54.1 78.2 +24.1 61.6 47.0 -14.6
LLaVA-1.5-7B 82.7 74.7 -8.0 93.2 133.4 +40.2
InternVL 158.3 93.4 -65.0 177.4 176.2 -1.2
6 Conclusion
Limitations
[…]
Computational Efficiency. Our implementation of PoCa involves more inferences to generate captions and prompting LLM for fusing the local and global captions. These multiple inference steps and the use of large models can lead to increased computational costs. This computational overhead may be a concern, especially in resource-constrained environments or when processing a large number of images. One potential solution is to finetune an image captioning model on the captions generated by PoCa. By doing so, the knowledge captured by PoCa can be distilled into the finetuned model, allowing for a single inference pass during deployment, while still benefiting from the enhanced caption quality achieved by PoCa. Similar approach of knowledge distillation has been adopted in other literature, such as Chen et al. (2023a); Fan et al. (2023); Chen et al. (2024).
Appendix
Prompt for Merging Caption Pyramid
System Message:
Input:
$\quad\bullet\quad$ You will receive a global caption describing an image.
$\quad\bullet\quad$ Additionally, you will have access to local captions generated for specific patches within the image.
$\quad\bullet\quad$ Both global and local captions may contain noise or errors.
Task Objective:
$\quad\bullet\quad$ Your goal is to create a merged global caption that combines relevant information from both sources.
$\quad\bullet\quad$ The merged caption should be no longer than the original ones.
$\quad\bullet\quad$ You only give the merged caption as output, without any additional information.
$\quad\bullet\quad$ Do NOT give any explaination or notes on how you generate this caption.
Guidelines:
$\quad\bullet\quad$ Combine Information: Extract key details from both global and local captions.
$\quad\bullet\quad$ Filter Noise: Remove non-sense content, inaccuracies, and irrelevant information.
$\quad\bullet\quad$ Prioritize Visual Details: Highlight essential visual elements instead of feeling or atmosphere
$\quad\bullet\quad$ Be Concise: Use as few words as possible while maintaining coherence and clarity.
$\quad\bullet\quad$ Ensure Coherence: Arrange the merged information logically.
Remember, your output should be a high-quality caption that is concise, informative, and coherent!
User:
### Global Caption: {global caption}
### Top-left: {top-left caption}
### Bottom-left: {bottom-left caption}
### Top-right: {top-right caption}
### Bottom-left: {bottom-left caption}
Assistant Generation Prefix:
Here’s the merged caption:
Appendix A Related Work
A.1 Image Captioning
Image captioning lies at the intersection of computer vision and natural language processing, requiring both accurate visual recognition and coherent language generation abilities. In Section 2 we formally defined the objectives for this task, but it is worth noting that current efforts do not explicitly optimize that objective. The primary challenge is the difficulty of back-propagation through the discrete textual space $\mathcal{D}_{\mathrm{caption}}$, and efforts addressing this challenge involve adopting reinforcement learning Yu et al. (2022b); Feng et al. (2019) or aligning continuous latent spaces with language spaces Liu et al. (2023a); Yu et al. (2023). However, these approaches suffer from training instability and less satisfactory language coherence.
Most current methods rely on a surrogate methodology, where they use human annotators $f_{\mathrm{human}}$ to write captions and train models to imitate those captions. The underlying assumption is that human-written captions optimize the objective $J(\mathrm{human})$, which is achieved by providing instructions to crowd-sourced caption annotators. For example, the instructions for MS COCO Caption annotation Chen et al. (2015) include “Describe all the important parts of the scene” and “Do not describe unimportant details”, which are respectively connected to the information sufficiency term and minimal redundancy term in our objective.
An important trend in the image captioning field is the increasing focus on the comprehensiveness of image captions. As mentioned earlier, this represents a decreased length penalty (smaller weight $\beta$ for the minimal redundancy term) and more emphasis on the information sufficiency term. In recent years, there has been an increasing number of high-quality detailed captioning datasets for this target, such as human-annotated image paragraph captioning Krause et al. (2017), Densely Captioned Images (DCI) Urbanek et al. (2023), and pseudo-labeled datasets, including LLaVA-Detailed-Captions Liu et al. (2023c), ShareGPT4V Chen et al. (2023a), AS-1B Wang et al. (2024), etc. However, detailed caption annotation is much more expensive than previous single-sentence annotation, while automated caption labeling exhibits a high risk of hallucination.
A.2 Vision-Language Learning in the Era of Large Language Models
Various methods have been explored for enabling vision-language learning in LLMs. One line of work focusing on vision-language alignment during pretraining Dosovitskiy et al. (2020); Chen et al. (2020); Zhang et al. (2021); Wang et al. (2022c, b); Yu et al. (2022a); Wang et al. (2022a); Radford et al. (2021); Li et al. (2022, 2023), allowing the model to jointly learn a shared vision-language latent space. The other line of work, improve the vision-language training efficiency by aligning the vision representation into the language space of LLMs by only training the visual encoder module or a vision-language projection matrix Tsimpoukelli et al. (2021); Alayrac et al. (2022); Liu et al. (2023c); Sun et al. (2023); Bai et al. (2023); Zhang et al. (2024). These two lines of works enable vision-language alignment, enabling various joint vision and language modalities prompting methods such as robot manipulation prompting Jiang et al. (2023); Wake et al. (2023) and multimodal in-context learning Zhao et al. (2024); Zhang et al. (2023).
Unlike the other two directions, another line of work exploits the reasoning and planning ability of LLMs allowing zero-shot multimodal vision-language inference by extracting the information from the vision modality into a textual description and performing inference through frozen LLMs Yang et al. (2022); Hu et al. (2022); Lin et al. (2022); Gui et al. (2022). Recent works in this direction showcase remarkable VQA ability through answer heuristics generation Shao et al. (2023) and enabling object tagging and image editing through visual programming Gupta and Kembhavi (2023). Inspired by this line of work, our work introduces a zero-shot hierarchical image captioning approach that relies on the reasoning ability of LLMs to aggregate information from local and global captions.
Appendix B VQA-based Caption Evaluation
@@
In this paper, we adopt the methodology of employing text-only LLMs for VQA inference with image captioning input to evaluate caption quality. This section provides a detailed analysis of this approach. Using the instruction given in Table 1, we prompt different LLMs to generate answers and evaluate the accuracy based on exact matching and Natural Language Inference (NLI) based evaluation. The NLI evaluation classifies a pair of statements, "The answer to this question is {ground truth}" and "The answer to this question is {generated answer}", into entailment, neutral, and contradiction, where entailment outputs are regarded as successful. Compared to exact matching, NLI evaluation measures the correctness of answers at a semantic level and can tolerate low-level differences.
@@
Appendix C PoCa Examples
We show some example of PoCa caption merging, where the images are sourced from the test split of the image paragraph captioning dataset  [^1] https://cs.stanford.edu/people/ranjaykrishna/im2p/index.html licensed by CC BY 4.0.. Captions are generated by LLaVA-1.5-7B model and merged by Mixtral-8x7B LLM. It shows that local captions successfully captured some errors that are hallucinated or omitted by the global caption, while PoCa successfully aggregated such information (marked by green) and recompile it into a human-interpretable caption.